### 26 Feb 2024 v1

@@ 1 Introduction
To emphasize the need for sample-efficient learning, consider again the application of closed-loop AFC to vehicles. The current gold standard for accurate and fast simulations of vehicle aerodynamics are hybrid approaches like improved delayed detached eddy simulations (IDDES) [11]. Computing converged mean force coefficients for the DrivAer test case at $Re=768000$, required approximately two days on 700 CPU cores in 2016 [11]. It is conceivable that roughly five sequential IDDES can be performed per day because it is not necessary to obtain converged statistics in every run. Moreover, hardware and scalability have improved significantly over the past years. Assuming that 10 simulations are performed in parallel per episode to generate more data, each simulation being run on 1000 CPU cores, the computational cost for 100 episodes is approximately 5M CPU hours, and the training lasts 20 days. Cloud providers charge 0.01-0.05 euros per CPU hour, which puts a price tag of 0.5M-2.5M euros on a single training at a single flow condition. Additional hyperparameter optimization, sensor/actuator variants, and an extension to multiple flow conditions can easily increase the cost by a factor of 10-100. This cost, combined with a turnaround time of 20 days, is far from practical for the majority of potential users.
@@
Several remedies exist to reduce the computational cost and turnaround time, e.g., by exploiting invariances [12, 13] or by pre-training on simulations with very coarse meshes [14]. While effective, the applicability of the aforementioned techniques strongly depends on the control problem, i.e., the problem must exhibit invariances, and the mesh coarsening must not change the flow characteristics. A more general approach to improve data efficiency is model-based DRL (MBDRL) [15]. The main idea is to substitute the expensive simulation-based environment with one or more surrogate models. The surrogate models are optimized based on data coming from the high-fidelity environment. Additional data to optimize the control law can be created with little effort by querying the optimized surrogate model. A plethora of MBDRL algorithms exist that differ in the type of surrogate model and how the control law is derived from the models. Moerland et al. [15] provide an extensive overview.
@@
Key challenges in MBDRL are i) the efficient creation of accurate surrogate models and ii) dealing with the presence of model error. Due to their flexibility, neural networks are a common choice for building data-driven surrogate models. However, the models must be created on the fly based on data that is unavailable for prior tests, so the optimization pipeline must be robust and efficient. Moreover, the models are auto-regressive. Hence, long-term stability is a persistent challenge. If the prediction accuracy is not sufficiently high, querying the models repeatedly leads to vanishing or exploding solutions. Finally, even if the model creation workflow is robust and efficient, there is always epistemic uncertainty due to the alternating iterative updates of environment models and control law. With each update of the control law, previously unknown actuations and states will emerge, and consequently, the environment models will fail at some point. Estimating if the surrogate models are still reliable or if they should be updated with high-fidelity data from the true environment is essential to make MBDRL robust.
@@
In this contribution, we present a modified version of the model ensemble trust region policy optimization (METRPO) [16] algorithm and demonstrate the benefits of MBDRL for fluid-mechanical applications. Specifically, we compare model-free (MF) and model-based (MB) learning on the flow past a circular cylinder [17, 18] and the fluidic pinball configuration [19, 9]. Moreover, we provide a complimentary code repository with instructions to reproduce the numerical experiments and data analyses presented hereafter at https://github.com/JanisGeise/MB_DRL_for_accelerated_learning_from_CFD.
@@ 2 Theory
This section starts with a very short introduction to essential RL concepts and then describes the MF algorithm employed to solve the flow control problems, namely proximal policy optimization (PPO). Thereafter, we explain the creation of model ensembles, which emulate the simulation outputs, and how these ensembles are employed in an MBDRL variant of PPO.
@@ 2.1 Reinforcement learning
There are two main entities in RL: the agent and the environment. The agent contains the control logic, while the environment represents the simulation or experiment that shall be controlled. The environment at a discrete time step $n\in\mathbb{N}$ is characterized by a state $S_{n}\in\mathcal{S}$, where $\mathcal{S}$ is the space of all possible states. Formally, there is a difference between the system’s state and a partial observation of the state, since the full state is often inaccessible. Nonetheless, we follow standard practice and refer to partial observations and full states alike simply as states. In fluid mechanics, $S_{n}$ could represent instantaneous pressure or velocity values at one or more sensor locations in the region of interest. The agent can manipulate the state via an action $A_{n}\in\mathcal{A}$, where $\mathcal{A}$ is the space of all possible actions, upon which the environment transitions to a new state $S_{n+1}$. Common actions (means of actuation) for flows are suction, blowing, rotation, or heating. The control objective in RL is expressed by the reward $R_{n+1}\in\mathcal{R}$, where $\mathcal{R}$ is the space of all possible rewards, and is assessed based on the transition from $S_{n}$ to $S_{n+1}$. For example, if the control objective is drag reduction, the reward would be defined such that large drag forces yield low rewards while small drag forces yield high rewards. The combination of $S_{n}$, $A_{n}$, $R_{n+1}$, and $S_{n+1}$ is called an experience tuple [20]. The aggregation of experience tuples of $N$ interactions between agent and environment forms a so-called trajectory [20], $\tau=\left[(S_{0},A_{0},R_{1},S_{1}),\ldots,(S_{N-1},A_{N-1},R_{N},S_{N})\right]$. Trajectories form the basis for optimizing the policy $\pi(A_{n}=a|S_{n}=s)$ (the control law), which predicts the probability of taking some action $a$ given that the current state is $s$. Optimality in RL is defined in terms of cumulative rewards, the so-called return $G_{n}=\sum_{i=n+1}^{N}\gamma^{i}R_{i}$, because the system’s response to an action may be delayed. The discount factor $\gamma\in(0,1]$ introduces a notion of urgency, i.e., the same reward counts more if achieved early in a trajectory. Given the same policy $\pi$ and the same state $S_{n}$, the return $G_{n}$ might vary between trajectories. Incomplete knowledge of the state is one common source of uncertainty. To consider uncertainty in the optimization, RL ultimately aims to find the policy that maximizes the expected return $v_{\pi}(s)$ under policy $\pi$ over all possible states $s\in\mathcal{S}$ [20]:
@@ 2.2 Policy learning
One sequence of generating trajectories, updating the value network according to (2), and updating the policy according to (6) is called an episode. Since the optimization starts with randomly initialized value and policy networks, multiple episodes are required to find the optimal policy.
@@ 2.3 Model learning
where $D$ is a set of feature-label-pairs constructed from high-fidelity trajectories, and $|D|$ is the overall number of feature-label-pairs. A single trajectory of length $N$ yields $|D|=N-(d+1)$ such pairs. To obtain the model weights based on (9), we employ state-of-the-art deep learning techniques, e.g., normalization of features and labels, layer normalization, batch training, learning rate schedule, and early stopping. More details about the model training are provided in appendix A.
@@ 3 Results
We demonstrate the MEPPO algorithm on two flow configurations, namely a rotating cylinder in channel flow and the fluidic pinball. The numerical setups are described in section 3.1. In section 3.2, we compare MF and MB trainings and investigate the influence of ensemble size $N_{\mathrm{m}}$ and switching criterion $N_{\mathrm{thr}}$. Section 3.3 presents a short discussion of the optimal policies found for each test case. The numerical simulations are performed with the OpenFOAM-v2206 toolbox. The orchestration of simulations and the DRL logic are implemented in the drlFoam package. Further implementation details are available in the complementary code repository.
@@ 3.1.1 Cylinder flow
The setup used in the present study is depicted in figure 1. The inlet velocity profile is parabolic [24], while the velocity vector at the upper and lower domain boundary is zero. At the outlet boundary, the velocity gradient is set to zero. For the pressure, a fixed reference value is applied at the outlet, while the gradient is set to zero on all other boundaries. The Reynolds number based on the mean inlet velocity $U_{\mathrm{in}}$, the cylinder diameter $d$ and the kinematic viscosity $\nu$ is $Re=U_{\mathrm{in}}d/\nu=100$.
@@
where the reference area is $A_{\mathrm{ref}}=dz$. Note that the setup is 2D, but 2D setups in OpenFOAM always have a single cell layer of depth $z$ in the third direction. The instantaneous reward at time step $n$ is computed as [18]:
@@ 3.1.2 Fluidic pinball
The setup is depicted in figure 2. To reduce the initial transient simulation phase, we apply a step function velocity profile at the inlet with $\varepsilon=0.01U_{\mathrm{in}}$. The same velocity vector is set on the lower and upper domain boundaries, i.e., $1.01U_{\mathrm{in}}$ and $0.99U_{\mathrm{in}}$, respectively. The velocity gradient at the outlet boundary is set to zero. Similar to the cylinder setup, the pressure is fixed at the outlet, and the pressure gradient on all other boundaries is set to zero.
@@
where the reference area is the projected area of all three cylinders in $x$-direction, namely $A_{\mathrm{ref}}=2.5dz$. Definition (13) ensures that the sum over all cylinders recovers the correct total force coefficients:
@@ 3.2.2 Cylinder flow
The markers in the MB training in figure 3 indicate CFD trajectories. For the MB training with one model, we switch back to simulation-based sampling every fourth episode, so $75\%$ of all episodes use the environment model. The same ratio results in the MEPPO training with $N_{\mathrm{m}}=5$ and $N_{\mathrm{thr}}=3$. Interestingly, the automated switching between CFD and model sampling leads to a relatively regular alternation. The same number of models with a lowered threshold value of $N_{\mathrm{thr}}=3$ reduces the amount of CFD episodes to $15\%$, notably without performance degradation. The training with $N_{\mathrm{m}}=10$ and $N_{\mathrm{thr}}=5$ also employs the model ensemble in $75\%$ of the episodes. Reducing the threshold value in the latter case to $N_{\mathrm{thr}}=3$ reduces the number of CFD episodes to $6\%$. However, the optimization becomes significantly less stable and does not reach optimal control performance. Note that the amount of training data per model decreases as the number of models increases. Therefore, increasing the number of sampled trajectories per episode or allowing overlapping training data could potentially stabilize trainings with a lowered threshold value. The increased amount of training data when employing only a single model could explain the quick convergence of the corresponding training after approximately 100 episodes.
Figure 4: Cylinder flow: composition of the total MB training time $T_{\mathrm{MB}}$ normalized with the MF training time $T_{\mathrm{MF}}$.
@@ 3.2.3 Fluidic pinball
Figure 6: Fluidic pinball: episode-wise mean reward $R$ for different training configurations; the shaded area encloses one standard deviation below and above the mean; mean and standard deviation are computed over all trajectories of seeds; for the MB training, the markers indicate CFD-based trajectory sampling.
@@
Figure 7: Fluidic pinball: composition of the total MB training time $T_{\mathrm{MB}}$ normalized with the MF training time $T_{\mathrm{MF}}$.
@@ 4 Conclusion
DRL-based AFC is a promising approach to enable smart control technology. A current limitation of simulation-based DRL is the high computational cost and turnaround time associated with the optimization. The idea of MBDRL is the creation of low-cost environment models that partially replace the costly sampling of high-fidelity trajectories from simulations. On two common benchmark AFC problems, we demonstrate that the MEPPO adopted here can tremendously reduce the training cost and time while achieving optimal control performance. The relative reduction of training time increases with the cost of the CFD simulation. Therefore, we expect the MEPPO algorithm, or variants thereof, to be a key enabler in performing DRL-based AFC on realistic, 3D simulations at an industrial scale.
There are a number of promising options to reduce the training cost even further. Creating accurate auto-regressive environment models on the fly posed a significant challenge. Automating the model creation would be extremely beneficial in the application of MBDRL to new problems. The more accurate the models can extrapolate, the more savings are possible. More advanced recurrent network architectures or transformers [27] might be capable of achieving higher accuracy than the simple, fully connected networks employed here. However, such advanced networks are typically also harder to train. Alternatively, more straightforward approaches to creating reduced-order models might be worthwhile to explore, too. For example, dynamic mode decomposition with control [28] could be a drop-in replacement for the models employed here.

### 10 Apr 2024 v2

@@ 1 Introduction
To emphasize the need for sample-efficient learning, consider again the application of DRL-based closed-loop AFC to vehicles. The current gold standard for accurate and fast simulations of vehicle aerodynamics are hybrid approaches like improved delayed detached eddy simulations (IDDES) [11]. Computing converged mean force coefficients for the DrivAer test case at $Re=768000$, required approximately two days on 700 CPU cores in 2016 [11]. It is conceivable that roughly five sequential IDDES can be performed per day because it is not necessary to obtain converged statistics in every run. Moreover, hardware and scalability have improved significantly over the past years. Assuming that 10 simulations are performed in parallel per episode to generate more data, each simulation being run on 1000 CPU cores, the computational cost for 100 episodes is approximately 5M CPU hours, and the training lasts 20 days. Cloud providers charge 0.01-0.05 euros per CPU hour, which puts a price tag of 0.5M-2.5M euros on a single training at a single flow condition. Additional hyperparameter optimization, sensor/actuator variants, and an extension to multiple flow conditions can easily increase the cost by a factor of 10-100. This cost, combined with a turnaround time of 20 days, is far from practical for the majority of potential users.
@@
Several remedies exist to reduce the computational cost and turnaround time, e.g., by exploiting invariances [12, 13] or by pre-training on simulations with very coarse meshes [14]. While effective, the applicability of the aforementioned techniques strongly depends on the control problem, i.e., the problem must exhibit invariances, and the mesh coarsening must not change the flow characteristics. A more general approach to improve data efficiency is model-based DRL (MBDRL) [15]. The main idea is to substitute the expensive simulation-based environment with one or more surrogate models. The surrogate models are optimized based on data coming from the high-fidelity environment. Additional data to optimize the control law can be created with little effort by querying the optimized surrogate model. A plethora of MBDRL algorithms exist that differ in the type of surrogate model and how the control law is derived from the model. Moerland et al. [15] provide an extensive overview.
@@
Key challenges in MBDRL are i) the efficient creation of accurate surrogate models and ii) dealing with the presence of model error. Due to their flexibility, neural networks are a common choice for building data-driven surrogate models. However, the models must be created on the fly based on data that is unavailable for prior tests, so the optimization pipeline must be robust and efficient. Moreover, the models are auto-regressive. Hence, long-term stability is a persistent challenge. If the prediction accuracy is not sufficiently high, querying the models repeatedly leads to vanishing or exploding solutions. Finally, even if the model creation workflow is robust and efficient, there is always epistemic uncertainty due to the alternating iterative updates of environment models and control law. With each update of the control law, previously unknown actuations and states will emerge, and consequently, the environment models will eventually become too inaccurate. Estimating whether the surrogate models are still reliable or should be updated with high-fidelity data from the true environment is essential to making MBDRL robust.
@@
In this contribution, we present a modified version of the model ensemble trust region policy optimization (METRPO) [16] algorithm and demonstrate the benefits of MBDRL for fluid-mechanical applications. Specifically, we compare model-free (MF) and model-based (MB) learning on the flow past a circular cylinder [17, 18] and the fluidic pinball configuration [19, 9]. The appendix describes various implementation details and serves as a reference for the abbreviations used throughout the text. Moreover, we provide a complimentary code repository with instructions to reproduce the numerical experiments and data analyses presented hereafter [20]. A persistent record of all scientific results, including a snapshot of the code repository, is publicly accessible [21].
@@ 2 Theory
This section starts with a very short introduction to essential RL concepts and then describes the MF algorithm, proximal policy optimization (PPO), employed to solve the flow control problems. Thereafter, we explain the creation of model ensembles, which emulate the simulation outputs, and how these ensembles are employed in an MBDRL variant of PPO. Figure 1 shows how the different algorithmic components are connected within one episode of model-based PPO.
Figure 1: High-level overview of one model-based PPO episode.
@@ 2.1 Reinforcement learning
There are two main entities in RL: the agent and the environment. The agent contains the control logic, while the environment represents the simulation or experiment that shall be controlled. The environment at a discrete time step $n\in\mathbb{N}$ is characterized by a state $S_{n}\in\mathcal{S}$, where $\mathcal{S}$ is the space of all possible states. Formally, there is a difference between the system’s state and a partial observation of the state, since the full state is often inaccessible. Nonetheless, we follow standard practice and refer to partial observations and full states alike simply as states. In fluid mechanics, $S_{n}$ could represent instantaneous pressure or velocity values at one or more sensor locations in the region of interest. The agent can manipulate the state via an action $A_{n}\in\mathcal{A}$, where $\mathcal{A}$ is the space of all possible actions, upon which the environment transitions to a new state $S_{n+1}$. Common actions (means of actuation) for flows are suction, blowing, rotation, or heating. The control objective in RL is expressed by the reward $R_{n+1}\in\mathcal{R}$, where $\mathcal{R}$ is the space of all possible rewards, and is assessed based on the transition from $S_{n}$ to $S_{n+1}$. For example, if the control objective is drag reduction, the reward would be defined such that large drag forces yield low rewards while small drag forces yield high rewards. The combination of $S_{n}$, $A_{n}$, $R_{n+1}$, and $S_{n+1}$ is called an experience tuple [22]. The aggregation of experience tuples over $N$ interactions between agent and environment forms a so-called trajectory [22], $\tau=\left[(S_{0},A_{0},R_{1},S_{1}),\ldots,(S_{N-1},A_{N-1},R_{N},S_{N})\right]$. Trajectories form the basis for optimizing the policy $\pi(A_{n}=a|S_{n}=s)$ (the control law), which predicts the probability of taking some action $a$ given that the current state is $s$. Optimality in RL is defined in terms of cumulative rewards, the so-called return $G_{n}=\sum_{i=n+1}^{N}\gamma^{i}R_{i}$, because the system’s response to an action may be delayed. The discount factor $\gamma\in(0,1]$ introduces a notion of urgency, i.e., the same reward counts more if achieved early in a trajectory. Given the same policy $\pi$ and the same state $S_{n}$, the return $G_{n}$ might vary between trajectories. Incomplete knowledge of the state is one common source of uncertainty. To consider uncertainty in the optimization, RL ultimately aims to find the policy that maximizes the expected return $v_{\pi}(s)$ under policy $\pi$ over all possible states $s\in\mathcal{S}$ [22]:
@@ 2.2 Policy learning
One sequence of generating trajectories, updating the value network according to equation (2), and updating the policy according to equation (6) is called an episode. Since the optimization starts with randomly initialized value and policy networks, multiple episodes are required to find the optimal policy.
@@ 2.3 Model learning
where $D$ is a set of feature-label-pairs constructed from high-fidelity trajectories, and $|D|$ is the overall number of feature-label-pairs. A single trajectory of length $N$ yields $|D|=N-(d+1)$ such pairs. To obtain the model weights based on equation (9), we employ state-of-the-art deep learning techniques, e.g., normalization of features and labels, layer normalization, batch training, learning rate schedule, and early stopping. More details about the model training are provided in appendix A.
@@ 3 Results
We demonstrate the MEPPO algorithm on two flow configurations, namely a rotating cylinder in channel flow and the fluidic pinball. The numerical setups are described in section 3.1. In section 3.2, we compare MF and MB trainings and investigate the influence of ensemble size $N_{\mathrm{m}}$ and switching criterion $N_{\mathrm{thr}}$. Section 3.3 presents a short discussion of the optimal policies found for each test case. The numerical simulations are performed with the OpenFOAM-v2206 toolbox [26]. The orchestration of simulations and the DRL logic are implemented in the drlFoam package [27]. Further implementation details are available in the complementary code repository [20].
@@ 3.1.1 Cylinder flow
The setup used in the present study is depicted in figure 2. The inlet velocity profile is parabolic [28], while the velocity vector at the upper and lower domain boundary is zero. At the outlet boundary, the velocity gradient is set to zero. For the pressure, a fixed reference value is applied at the outlet, while the gradient is set to zero on all other boundaries. The Reynolds number based on the mean inlet velocity $U_{\mathrm{in}}$, the cylinder diameter $d$ and the kinematic viscosity $\nu$ is $Re=U_{\mathrm{in}}d/\nu=100$. To solve the incompressible Navier Stokes equations, we employ the pimpleFoam solver with residual control. The pressure-velocity coupling is stopped once the initial residuals for pressure and momentum equations drop below $10^{-4}$. Since the flow is laminar, no additional turbulence modeling is necessary. The mesh consists of approximately $5.3\times 10^{3}$ hexahedral cells and is created using the blockMesh utility. For completeness, we note that the mesh has one cell layer in the third spatial direction of depth $\Delta z=0.1d$, even though the simulation is 2D. The extension in the third direction is a technical requirement of 2D simulations in OpenFOAM. To discretize convective and diffusive fluxes, we employ pure linear interpolation. An implicit first-order Euler method is used for the temporal discretization.
@@
where the reference area is $A_{\mathrm{ref}}=d\Delta z$. The instantaneous reward at time step $n$ is computed as [18]:
@@ 3.1.2 Fluidic pinball
The setup is depicted in figure 3. To reduce the initial transient simulation phase, we apply a step function velocity profile at the inlet with $\varepsilon=0.01U_{\mathrm{in}}$. A constant velocity vector is set on the lower and upper domain boundaries, i.e., $1.01U_{\mathrm{in}}$ and $0.99U_{\mathrm{in}}$, respectively. The velocity gradient at the outlet boundary is set to zero. Similar to the cylinder setup, the pressure is fixed at the outlet, and the pressure gradient on all other boundaries is set to zero. The remaining details of the setup are largely the same as those presented in section 3.1.1. Due to the more complex geometry, the mesh consists of approximately $5.3\times 10^{4}$ control volumes. For the convective term of the momentum equation, we employ a limitedLinear scheme with a coefficient of $1.0$.
@@
where the reference area is the projected area of all three cylinders in $x$-direction, namely $A_{\mathrm{ref}}=2.5d\Delta z$. Definition (13) ensures that the sum over all cylinders recovers the correct total force coefficients:
@@ 3.2.2 Cylinder flow
The markers in the MB training in figure 4 indicate CFD trajectories. For the MB training with one model, we switch back to simulation-based sampling every fourth episode, so $75\%$ of all episodes use the environment model. The same ratio results in the MEPPO training with $N_{\mathrm{m}}=5$ and $N_{\mathrm{thr}}=3$. Interestingly, the automated switching between CFD and model sampling leads to a relatively regular alternation. The same number of models with a lowered threshold value of $N_{\mathrm{thr}}=2$ reduces the amount of CFD episodes to $15\%$, notably without performance degradation. The training with $N_{\mathrm{m}}=10$ and $N_{\mathrm{thr}}=5$ also employs the model ensemble in $75\%$ of the episodes. Reducing the threshold value in the latter case to $N_{\mathrm{thr}}=3$ reduces the number of CFD episodes to $6\%$. However, the optimization becomes significantly less stable and does not reach optimal control performance. Note that the amount of training data per model decreases as the number of models increases. Therefore, increasing the number of sampled trajectories per episode or allowing overlapping training data could potentially stabilize trainings with a lowered threshold value. The increased amount of training data when employing only a single model could explain the quick convergence of the corresponding training after approximately 100 episodes.
Figure 5: Cylinder flow: composition of the total MB training time $T_{\mathrm{MB}}$ normalized with the MF training time $T_{\mathrm{MF}}\approx 14h$.
@@ 3.2.3 Fluidic pinball
Figure 7: Fluidic pinball: episode-wise mean reward $R$ for different training configurations; the shaded area encloses one standard deviation below and above the mean; mean and standard deviation are computed over all trajectories; for the MB training, the markers indicate CFD-based trajectory sampling.
@@
Figure 8: Fluidic pinball: composition of the total MB training time $T_{\mathrm{MB}}$ normalized with the MF training time $T_{\mathrm{MF}}\approx 130h$.
@@ 4 Conclusion
DRL-based AFC is a promising approach to enable smart control technology. A current limitation of simulation-based DRL is the high computational cost and turnaround time associated with the optimization. The idea of MBDRL is the creation of low-cost environment models that partially replace the costly sampling of high-fidelity trajectories from simulations. On two common benchmark AFC problems, we demonstrate that the MEPPO algorithm adopted here can tremendously reduce the training cost and time while achieving optimal control performance. The relative reduction of training time increases with the cost of the CFD simulation. Therefore, we expect the MEPPO algorithm, or variants thereof, to be a key enabler in performing DRL-based AFC on realistic, 3D simulations at an industrial scale. Of course, further tests on more complex flow configurations are required to consolidate this hypothesis.
There are a number of promising options to reduce the training cost even further. Creating accurate auto-regressive environment models on the fly posed a significant challenge. Automating the model creation, e.g., by means of Bayesian hyperparameter optimization, could be beneficial for the models’ accuracy and adaptability to new control problems. The more accurate the models can extrapolate, the more savings are possible. More advanced recurrent network architectures or transformers [31] might be capable of achieving higher accuracy than the simple, fully connected networks employed here. However, such advanced networks are typically also harder to train. Alternatively, more straightforward approaches to creating reduced-order models might be worthwhile to explore, too. For example, dynamic mode decomposition with control [32] could be a drop-in replacement for the models employed here.
@@
Ethical approval
No experiments on humans or animals have been conducted.
Funding
This work has been funded by the German Research Foundation (DFG) under the grant WE 6948/1-1.
Availability of data and materials
The full research data and instructions to reproduce the numerical experiments and analyses are available at:
https://doi.org/10.23728/b2share.85ab8f3f68724372b83babbdaca85910